OpenStudy (anonymous):

If \(\sec \theta - \tan \theta = \frac{1}{2}\), then \(\theta\) lies in which quadrant ?

OpenStudy (anonymous):

sec^2x-tan^2x=1

OpenStudy (anonymous):

sec theta=cos(90-theta)

OpenStudy (anonymous):

cot(90-theta)=tan theta

OpenStudy (lochana):

not clear , so do we need to proof sec^2x-tan^2x=1,sec theta=cos(90-theta),cot(90-theta)=tan theta

OpenStudy (anonymous):

(secx-tanx)2=(1/2)^2

OpenStudy (anonymous):

lies in which quadrant

OpenStudy (anonymous):

@ikram002p

ganeshie8 (ganeshie8):

I have modified the question, see if it looks okay now :)

ganeshie8 (ganeshie8):

refresh the page to see the changes

OpenStudy (anonymous):

yes .but it is not clear to me its not loaded fully

ganeshie8 (ganeshie8):

see if this works : \[\large \sec^2\theta - \tan^2\theta = 1\]

ganeshie8 (ganeshie8):

\[\large (\sec\theta + \tan\theta)(\sec\theta - \tan\theta) = 1\] \[\large (\sec\theta + \tan\theta)(\frac{1}{2}) = 1\] \[\large \sec\theta + \tan\theta = 2 \tag{2}\]

ganeshie8 (ganeshie8):

add this to the given equation : \[\large \begin{array} \\ \sec\theta - \tan\theta &= \frac{1}{2} \\ \sec\theta + \tan\theta &= 2 \\ \end{array} \]

ganeshie8 (ganeshie8):

add them gives you : \(\large 2\sec \theta = 5/2\) or \(\large \cos \theta = 4/5\)

ganeshie8 (ganeshie8):

subtracting them gives you : \(\large 2\tan \theta = 3/2\) or \(\large \tan \theta = 3/4\)

ganeshie8 (ganeshie8):

so both tan and cos are positive, in which Quadrant are they both positive?

OpenStudy (anonymous):

90 first quadrant. i dont know quadrants

OpenStudy (anonymous):